OpenStudy (anonymous):

For what value a is the tangent line to the graph f(x) = -x^2 + x +c at the point (a,f(a)) parallel to the line y= -x+3

OpenStudy (sirm3d):

you have a tangent line (whose slope is determined by the derivative) parallel to the given line y=-x+3 whose slope is...

OpenStudy (anonymous):

i found the derivative of -x^2 + x + which is -2x + 1

OpenStudy (anonymous):

and i equate -2x + 1 = -x +3

OpenStudy (anonymous):

and is wrong ^^ lol

OpenStudy (sirm3d):

for some x, the slope of the tangent line is equal to the slope of the line y = -x + 3 (whose slope is -1)

OpenStudy (sirm3d):

solve x in -2x + 1 (slope of tangent line) = -1 (slope of the given line)

OpenStudy (anonymous):

but the given answer says +1

OpenStudy (anonymous):

oh yes

OpenStudy (anonymous):

slope is -1 , a = 1

OpenStudy (anonymous):

but what's the point of -x^2+x+c

OpenStudy (sirm3d):

the is the curve where one of its tangent lines (whose slope is determined by its derivative) is parallel to some given line, in this problem, y= -x + 3. if there's no curve, there's no point discussing tangent lines.

OpenStudy (anonymous):

well, there is no graph given so basically if the function f(x) has a tangent whose equation is given, the slope AT ONE POINT of the function is the same as y=-x+3 where as -2x+x defines the slope at all points then can't I equate -2x+x = -x+3 to find the point that they are equal?

OpenStudy (sirm3d):

-2x + 1 is THE slop of tangent line, while -x + 3 is a line. you can't equate the slope and a line. it must be slope = slope.

OpenStudy (anonymous):

then slope of the line -x+3 = -1 -2x+1 = -1 x= 1 oh.. lol

OpenStudy (sirm3d):

gotcha

OpenStudy (anonymous):

oh thanks!! and then another one find derivative of x^cos(x) i just forgot the formula to do that..

OpenStudy (sirm3d):

let y = x^cos x take the logarithm first of both sides, then use implicit differentiation

OpenStudy (anonymous):

can i use product rule

OpenStudy (anonymous):

i first did it and got -sinxlnx + cosx / x but i think the answer should include xcos^x in it

OpenStudy (sirm3d):

yes, after taking the logarithm of both sides.

OpenStudy (anonymous):

still dont get it lol

OpenStudy (anonymous):

after u log it , it becomes cos x ln x.. then where is the x ^cosx

OpenStudy (anonymous):

where u go @@

OpenStudy (sirm3d):

y=x^cos x ln y = ln (x^cos x) ln y = cos x ln x now differentiate both sides (1/y) dy/dx = cos x (1/x) - sin x ln x solve dy/dx

OpenStudy (anonymous):

oh ok thanks! and when i ask where u go, basically just to make sure u received the msg, so is not u wait for me , and i 'm waiting for u as well. so if u have anything to do, u can always just come back after

OpenStudy (sirm3d):

im having problems with my net connection.

OpenStudy (anonymous):

oh ok, and np, i will just wait :P

OpenStudy (anonymous):

and then #17, just sent u up there

OpenStudy (sirm3d):

\[\Large f'(x)=\begin{cases}1, & x>2\\-1, & x<2 \end{cases}\]

OpenStudy (sirm3d):

notice that in the derivative, the "=" disappeared in \(x \leq 2\)

OpenStudy (anonymous):

yes, that means it is not differentiable at x=2

OpenStudy (sirm3d):

right. differentiability is assured only on an open interval, not at an endpoint. in this case, x=2 is an endpoint.

OpenStudy (anonymous):

oh, answer is E not A because it includes x=2 being differentiable not B because of the same reason not C because x=2 is not differentiable Not D because everywhere is differentiable except x=2 so is E!

OpenStudy (anonymous):

and what does the 1, and -1 mean < i know is the slope, but what kind of info does it give?

OpenStudy (sirm3d):

f'(x) =1 , x > 2 is telling us that the derivative exists and its value is 1

OpenStudy (anonymous):

ohh oh yea, thanks! if u need to do my test, i feel like u are gonna finish it in 15mins ..>.>

OpenStudy (anonymous):

and #20 plz :D

OpenStudy (sirm3d):

what do you think?

OpenStudy (anonymous):

It was on my midterm and I just gave up lol

OpenStudy (anonymous):

for the dot product, i believe they need to be differentiable and i think there is a rule for that

OpenStudy (anonymous):

but I dont remember the rule, and no idea where to find it..and plus those options are confusing enough

OpenStudy (sirm3d):

it's (D). THe derivative of \(f \cdot g\) is \(f' \dot g + f \cdot g'\), PROVIDED \(f\) and \(g\) are differentiable. the required premises for computing the derivative is the differentiability of the factors.

OpenStudy (anonymous):

if a variable has a derivative, does it mean it is differentiable for sure

OpenStudy (sirm3d):

differentiable only on an open interval, NOT the closed interval.

OpenStudy (anonymous):

i actually dont understand what D is saying @@

OpenStudy (anonymous):

what is those "not guarantee" stuff... talking about open and closed intevals?

OpenStudy (sirm3d):

it says if f, g, f' and g' exists, then (1) (fg)' exists; and (2) we can use the formul (fg)' = f'g + fg'

OpenStudy (anonymous):

but it also says differentiability of f exist doesnt mean g exist but if f dot g exist, shouldnt g be differentiable as wlel

OpenStudy (anonymous):

like..how can u put something differnetiable with something that is not differentiable and get something differentiable @@

OpenStudy (sirm3d):

it only says you know tha (fg)' exists. you cannot use the formula (fg)' = fg' + gf' since you are not assured that both g' and f' exists.

OpenStudy (sirm3d):

to translate it, if you know that material A is magnetically attracted to B, can you be certain that A is a magnet?

OpenStudy (anonymous):

noo

OpenStudy (anonymous):

ahhh...much more clear :P

OpenStudy (anonymous):

thks for the analogy haha

OpenStudy (sirm3d):

(D) says how to compute (fg)' if you know f' and g'.

OpenStudy (anonymous):

yea , i get it when u explain. but I'm really weak in these theory/concept questions :/

OpenStudy (anonymous):

basically if a ques has too many words , i will just be lost..

OpenStudy (anonymous):

for x^2 - 2x^2 + 2x/ x^4 to find the derivative, is it better to break it in parts of just directly the difference rule

OpenStudy (sirm3d):

agree.

OpenStudy (anonymous):

so break it by parts?

OpenStudy (anonymous):

and then for f(x) = x^1/3 , why Newtons' method will fail with x1 = 1 is it because the point it taken near the end point, which is the absolute minimum? As I see Newton's method will fail if 1.) f'(x)=0 2.) The initial value is taken at a maximum or minumum point, 3.) The initial value is taken at a point of inflection, 4.) The initial value is taken near a maximum point and a minimum point, 5.) The initial value is taken near a point of inflection.

OpenStudy (sirm3d):

still thinking...

OpenStudy (anonymous):

oh ok